# Numerical efficiency of explicit time integrators for phase-field models

### v1

@@ 3.2 Temporal adaptivity
For conditionally stable schemes, often a time step close to the stability limit is chosen for reasons of computational efficiency, with the hope being that the temporal error introduced thus is much smaller than the spatial error. However, for unconditionally stable schemes some kind of adaptivity is necessary as to decide what step size to take. For this an estimate of the local error introduced by the time stepping process is useful. The SSP methods are extended with the embedded solution coefficients of Fekete et al. [2022] and thus they can estimate the error at little extra computational cost but with one extra vector being stored. For the STS methods, the error may be estimated following Sommeijer et al. [1998] requiring an extra RHS evaluation before a time step is accepted [^1] This evaluation might be re-used for the first stage of the next step if it’s accepted; for simplicity this wasn’t implemented and the additional RHS evaluations were counted. but without extra storage requirements. The error formula is then
$e^{n}_{u}$ $=\frac{1}{15}(12(u^{n+1}_{i}-u^{n}_{i})+6\Delta t(\operatorname{rhs}(u^{n}_{i})+\operatorname{rhs}(u^{n+1}_{i})))$ (33)
@@
In either case, once a local error estimate $e^{n}_{i}$ is available, the error of this step is calculated as
$e^{n}$ $=\left[\frac{1}{N}\sum_{i}|\frac{e^{n}_{i}}{r\max(u^{n}_{i},u^{n+1}_{i})+a}|^{2}\right]^{\frac{1}{2}}$ (34)
by computing a weighted $l^{2}$ norm of the error over all DoFs $i$. The weight is calculated with a relative tolerance $r$ and an absolute tolerance $a$ to define the level of acceptable error. These tolerances are defined per independent field, i.e. $r_{\phi},a_{\phi}$ for the phase-field and $r_{c},a_{c}$ for the concentration. However, since the components of the phase-field are tightly coupled via the simplex and sum constraints, with values to 1 and 0 occurring at the same location, the absolute phase-field tolerance $a_{\phi}$ will be varied within the work and its influence mentioned where appropriate..
If $e^{n}<1$, the step is accepted, otherwise rejected. In either case, a new time step is calculated following the PID-controller of Söderlind and Wang [2006] together with a bias $b\in[0.1,0.98]$ and the local order $P$ of the method
@@
with $k_{1}=1.25,k_{2}=0.5,k_{3}=-0.6,k_{4}=0.25,k_{5}=0$; note that the step index is now a subscript as not to be confused with the exponentiation. The bias $b$ starts at $0.9$ and makes the step estimate slightly more conservative; it is multiplied by itself on step rejection to make successive step rejections more unlikely, up to a value of $0.1$. On step acceptance, it is divided by $\sqrt{b}$ but limited to $0.98$. The time step factor $F$ is then smoothed to limit the overall step size changes. Terms for which there is no data yet are dropped from the above expression. We observed that the full controller may suggest increasing the time step upon rejection; in this case only $e^{-k_{1}/(P+1)}_{n}$ is used for estimating the new timestep, which is certain to decrease the time step.
@@
We note that reductions, such as required by the error estimator, are nondeterministic on GPUsShanmugavelu et al. [2024] and so the runtimes and RHS counts can slightly fluctuate when using an adaptive time step on GPUs. While this also applies to MPI parallel reductions, the number of non-associative operations is often far smaller and so can usually be ignored.
@@ 3.3 Error sources
The spatial discretization error for the Laplacian of the equilibrium phase-field profile eq. 7, away from the bulk, scales as
$e_{\Delta x}\propto\sum_{n=1}q_{n}\frac{\Delta x^{2n}}{W^{2+2n}}$ (38)
@@
and thus, in order for the error to vanish as $\Delta x,W\to 0$, the terms $\frac{\Delta x^{2n}}{W^{2+2n}}$ need to go to zero. For constant $W$, one obtains the usual error order if a reference solution using a sufficiently small $\Delta x$ is employed. However, this does not necessarily recover the sharp interface limit $W\to 0$ of the free-boundary problem. Thus, when doing grid convergence studies, $W$ is chosen as $W=3\Delta x^{m}$ with $m=0.4<\frac{1}{2}$ for which the above error goes to zero for any $n$. The prefactor is chosen such that at the coarsest resolution employed we have about 9 cells in the interface. The error w.r.t. the sharp interface solution does not necessarily follow a simple power law, since it is the combined error of discretization and from the diffuse interface.
4 Benchmark geometries and parametrization
@@
In this section the benchmark geometries will be introduced, along with how the error for each one will be estimated. The benchmarks are generally chosen such that an analytical solution for the equivalent free-boundary problem exists. This allows us to determine when spatial or temporal error dominate as well as what kind of tolerances can be appropriate. Furthermore, the parameters employed for the simulation will be detailed.
In most cases, the initial conditions of the phase-field are based on calculating the profile function eq. 7 with a suitably defined distance field $d$ as its argument. Thus only the distance field $d$ will be given for these cases. If a simulation includes eq. 16, then the initial concentration field is given by the local equilibrium eq. 10 unless mentioned otherwise. If not mentioned otherwise, all boundary conditions are periodic.
An important part of the phase-field theory is that the derived PDEs represent gradient flows of an energy functional. This ensures that the evolution minimizes that functional. Schemes ensuring this are generally referred to as energy stableGomez and Van Der Zee [2017], with many classical schemes not being unconditionally so. In lieu of proving (conditional) energy stability for schemes employed herein, two cases with analytical equilibrium state are considered. Even if the schemes should temporarily have increasing energy (as shown), it is deemed acceptable if the analytical equilibrium state is the convergence point of a grid refinement study and that at a practical resolution, the scheme approximates the equilibrium state closely as well. Note that all schemes employed here contain at least one (scaled) forward Euler step and hence would at best be conditionally energy stable. We further note that even energy stable schemes might not lead to the correct kineticsXu et al. [2019], so they should also be benchmarked before use.
@@
The standard set of parameters employed is shown in Table 3, which will be used unless mentioned otherwise; these do not represent any specific material. For equilibrium simulations, $D=100$ is used to speed up convergence.
@@
parameter value
$M$ 1
$\gamma$ 1
$W$ 2.5
$\Delta x$ 1
$D$ 1
$k_{\alpha}=k_{\beta}$ 500
$c_{0,\alpha}$ 0.02
$c_{0,\beta}$ 0.98
Finally, the values used for equilibrium determination in the benchmarks sections 4.1 and 4.2 may oscillate around a steady-state value. In order to account for this, an oscillation resistant-algorithm detailed in the supplementary material is employed.
4.1 Phase embedding
The first case is that of one phase being embedded in another, with curvature providing a driving force for shrinkage which is counteracted by chemical driving forces. Both phases are initialized with their flat equilibrium concentration values and thus there is a small driving force for diffusion in order to reach equilibrium in which there is a pressure difference, the Laplace pressure,
@@ 4.1 Phase embedding
between both phases, with the curvature $\kappa$. This pressure difference is achieved by a shift in chemical potential via mass transfer. In order to accommodate this mass transfer, the volume of the individual phases changes slightly. The curvature is determined from a shape assumption, e.g. for a circular embedding we have $\kappa=\frac{1}{r}$ with $r$ being the radius of the embedded phase. The pressure difference is equivalent to the difference in grand potentials $\Delta\psi$ of the bulk phases in equilibrium, as this is the driving force resisting capillary pressure. The bulk grand potentials are defined as the values the respective grand potentials take on at the average chemical potential of the domain. Hence the error is defined as
@@
We note that resolving a discretized profile with different interface widths tends to produce slightly different areas after integration, as there is some error associated with the interface width (see the supplementary material for details). This also affects the concentration in the domain, which in turn affects the equilibrium. In order to counteract this, the difference of the sharp-interface concentration $C=V_{e}c_{e}+V_{E}c_{E}$ (with the sharp interface volumes $V_{e}$ and $V_{E}$ of the embedding/embedded phases) to the concentration in the domain $\int_{V}cdV$ is used to shift all concentration values in the domain by their difference divided by the volume of the domain. This is also done for the next benchmark.
4.2 Double triple junction
The second case is a triple junction geometry, now also probing variable parameters and multiple phases interacting. One phase $\alpha$ is placed on the grain boundary between two grains of the phase $\beta$, as sketched in fig. 1 together with the employed boundary conditions and geometric quantities, forming two triple junctions. In contrast to Daubner et al. [2023], we want to test the equilibrium together with a chemical driving force, and so do not require Dirichlet boundary conditions to prevent curvature minimization from eliminating all interfaces. We note that this choice generally leads to much longer simulation times, as the equilibrium is now linked to the long-range diffusion field. The evolution should be such that the phase $\alpha$ obtains an equilibrium shape defined by the dihedral angle
@@ 4.2 Double triple junction
with its momentary form being calculatable with the geometric parameters $L$, $S$ following Bradley and Aaronson [1977] by assuming the shape is a vesica piscis:
@@
$\begin{bmatrix}1&-q_{12}&0&\cdots&0\\ 0&1&-q_{23}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 1&1&1&\cdots&1\end{bmatrix}\begin{bmatrix}\phi^{{}^{\prime}}_{1}\\ \phi^{{}^{\prime}}_{2}\\ \vdots\\ \phi^{{}^{\prime}}_{N}\\ \end{bmatrix}=\begin{bmatrix}0\\ 0\\ \vdots\\ 1-\phi_{\alpha}\end{bmatrix}$.
4.3 Single grain growth
A single grain embedded within another grain will shrink due to excess energy contained in the interface, given that there are no other driving forces preventing this. Assuming that curvature flow is the only active mechanism, one may write for the volume of the embedded grain

### v2

@@ 3.3 Temporal adaptivity
For conditionally stable schemes, often a time step close to the stability limit is chosen for reasons of computational efficiency, with the hope being that the temporal error introduced thus is much smaller than the spatial error. However, for unconditionally stable schemes some kind of adaptivity is necessary as to decide what step size to take. For this an estimate of the local error introduced by the time stepping process is useful. The SSP methods may be extended with an embedded method, i.e. a method requiring function evaluations at the same or a subset of stages, and which calculates a lower-order solution of the problem. The difference between the solutions may then be taken as a measure of the error, i.e. $e^{n}_{i}=u^{n}_{i}-z^{n}_{i}$ with $z$ being the embedded solution and $u$ being the normal solution. Hence embedded methods can estimate the error at little extra computational cost but with one extra vector being stored. The embedded methods for the SSP methods use the embedded solution coefficients suggested by Fekete2022 .
For the STS methods, the error may be estimated following Sommeijer1998 by a calculation involving the old and new solution. This requires an extra RHS evaluation before a time step is accepted [^1] This evaluation might be re-used for the first stage of the next step if it’s accepted; for simplicity this wasn’t implemented and the additional RHS evaluations were counted. but without extra storage requirements. The error formula is then
$e^{n}_{i}$ $=\frac{1}{15}(12(u^{n+1}_{i}-u^{n}_{i})+6\Delta t(\operatorname{rhs}(u^{n}_{i})+\operatorname{rhs}(u^{n+1}_{i})))$ (34)
@@
In either case, once a local error estimate $e^{n}_{i}$ is available, the error of this step is calculated asSommeijer1998
$e^{n}$ $=\left[\frac{1}{N}\sum_{i=0}^{N}|\frac{e^{n}_{i}}{r\max(u^{n}_{i},u^{n+1}_{i})+a}|^{2}\right]^{\frac{1}{2}}$ (35)
by computing a weighted $l^{2}$ norm of the error over all DoFs $i$. The weight is calculated with a relative tolerance $r$ and an absolute tolerance $a$ to define the level of acceptable error. These tolerances are defined per independent field, i.e. $r_{\phi},a_{\phi}$ for the phase-field and $r_{c},a_{c}$ for the concentration. One would usually choose the absolute tolerance at the level of error or noise one expects of a field, and so it may be seen as a threshold at which a value might as well be zero. For the concentration $c$ this lets us choose reasonable tolerances since we know the expected value range via thermodynamics and its smallest reasonable value doesn’t approach zero. For the phase-field however the entire interval $[0,1]$ represents possible values, with one cell often containing values close to both 11 and 0. Since the phase-field can get arbitrarily close to 0 and in fact can become 0, the choice of $a_{\phi}$ is not obvious. Hence it will be varied within the work and its influence mentioned where appropriate. While assigning different tolerances to the different entries of the phase-field is possible, each individual phase-field can range over $[0,1]$; only the sum is known to be 1. Future work may consider altering how the error is calculated for the phase-field.
If $e^{n}<1$, the step is accepted, otherwise rejected. In either case, a new time step is calculated following the PID-controller of Soderlind2006 together with a bias $b\in[0.1,0.98]$ and the order $P$ of the method
@@
with $k_{1}=1.25,k_{2}=0.5,k_{3}=-0.6,k_{4}=0.25,k_{5}=0$; note that the step index $n$ is now a subscript as not to be confused with the exponentiation. The bias $b$ starts at $0.9$ and makes the step estimate slightly more conservative; it is multiplied by itself on step rejection to make successive step rejections more unlikely, up to a value of $0.1$. On step acceptance, it is divided by $\sqrt{b}$ but limited to $0.98$. The time step factor $F$ is then smoothed to limit the overall step size changes. Terms for which there is no data yet are dropped from the above expression. We observed that the full controller may suggest increasing the time step upon rejection; in this case only $e^{-k_{1}/(P+1)}_{n}$ is used for estimating the new timestep, which is certain to decrease the time step.
@@
We note that reductions, such as required by the error estimator, are nondeterministic on GPUsShanmugavelu2024 and so the runtimes and RHS counts can slightly fluctuate when using an adaptive time step on GPUs. While this also applies to MPI parallel reductions, the number of non-associative operations is often far smaller and so can usually be ignored.
@@ 3.4 Error sources
A central finite difference discretization for the second derivative of a function $u(x)$ may be obtained by adding Taylor expansions for $u(x_{0}+\Delta x)$ and $u(x_{0}-\Delta x)$ and solving for the second derivative at $x_{0}$. This yields, noting that all odd derivatives cancel,
$\frac{d^{2}u}{dx^{2}}$ $\approx\frac{u(x_{0}+\Delta x)-2u(x_{0})+u(x_{0}-\Delta x)}{\Delta x^{2}}-2\sum_{n=2}^{\infty}\frac{\Delta x^{2n-2}}{(2n)!}\frac{d^{2n}u}{dx^{2n}}$ (39)
i.e. the standard second order finite difference, with the infinite sum at the end representing the error. Combine this with $\frac{d^{n}\phi}{dx^{n}}\propto\frac{1}{W^{n}}$ for the equilibrium profile eq. 8 away from the bulk via simple differentiation, shift indices, and we can see that the spatial discretization error scales as
$e_{\Delta x}\propto\sum_{n=1}^{\infty}q_{n}\frac{\Delta x^{2n}}{W^{2+2n}}$ (40)
@@
if we are sufficiently far from the bulk to ignore the case statement, with $q_{n}$ absorbing any remaining constants and factors depending on $n$ solely. Thus, in order for the error to vanish as $\Delta x,W\to 0$, the terms $\frac{\Delta x^{2n}}{W^{2+2n}}$ need to go to zero. For constant $W$, one obtains the usual error order if a reference solution using a sufficiently small $\Delta x$ is employed. However, this does not necessarily recover the sharp interface limit $W\to 0$ of the free-boundary problem. Thus, when doing grid convergence studies, $W$ is chosen as $W=3\Delta x^{m}$ with $m=0.4<\frac{1}{2}$ for which the above error goes to zero for any $n$. This scaling essentially means that the interface becomes increasingly well-resolved as the grid spacing $\Delta x$ is reduced. The prefactor is chosen such that at the coarsest resolution employed we have about 9 cells in the interface. Finally, note that the error w.r.t. the sharp interface solution does not necessarily follow a simple power law, since it is the combined error of discretization and from the diffuse interface.
4 Common conditions & parametrization
@@
In this section the common parameters and conditions will be explained, with the following section presenting single benchmarks and their solution together. The benchmarks are generally chosen such that an analytical solution for the equivalent free-boundary problem exists. This allows us to determine when spatial or temporal error dominate as well as what kind of tolerances can be appropriate. If not mentioned otherwise, all boundary conditions are periodic.
In most cases, the initial conditions of the phase-field are based on calculating the profile function eq. 8 with a suitably defined distance field $d$ as its argument. Thus only the distance field $d$ will be given for these cases. However, for some configurations the geometric description based solely on distance fields breaks down as there are multiple overlapping distance fields, whose summed phase-field value do not sum to 1; this is the case for sections 5.1.2 and 5.3. In this case we consistently proceed as follows: We pick one phase $\alpha$ to exactly follow the distance field $d_{\alpha}$ and thus we can fix its value $\phi_{\alpha}=\phi(d_{\alpha})$. Based on the distance fields of the remaining grains, we generate trial $\phi$ values whose ratios $q_{ij}=\frac{\phi_{i}}{\phi_{j}}$ we would like to preserve together with the sum constraint $1-\phi_{\alpha}=\sum_{\beta=0,\beta\neq\alpha}^{N-1}\phi_{\beta}$. This generates at first a nonlinear system for the new values $\phi^{{}^{\prime}}$ which satisfy these constraints. By multiplying all equations but the sum one by the unknown divisors we obtain a linear system, e.g. for $\alpha=0$:
$\begin{bmatrix}1&-q_{12}&0&\cdots&0\\ 0&1&-q_{23}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 1&1&1&\cdots&1\end{bmatrix}\begin{bmatrix}\phi^{{}^{\prime}}_{1}\\ \phi^{{}^{\prime}}_{2}\\ \vdots\\ \phi^{{}^{\prime}}_{N-1}\\ \end{bmatrix}=\begin{bmatrix}0\\ 0\\ \vdots\\ 1-\phi_{\alpha}\end{bmatrix}$. (41)
If a simulation includes eq. 17, then the initial concentration field is given by the local equilibrium eq. 11 unless mentioned otherwise. We note though that resolving a discretized profile with different interface widths tends to produce slightly different areas after integration, as there is some error associated with the interface width (see the supplementary material for details). In order to counteract this, the difference of the sharp-interface concentration $C=V_{e}c_{e}+V_{E}c_{E}$ (with the sharp interface volumes $V_{e}$ and $V_{E}$ of the embedding/embedded phases) to the concentration in the domain $\int_{V}cdV$ is used to shift all concentration values in the domain by their difference divided by the volume of the domain. This is only done for the equilibrium benchmarks sections 5.1.1 and 5.1.2 to ensure approximately the same equilibrium sizes are recovered under grid refinement.
@@
The values used for equilibrium determination in the benchmarks sections 5.1.1 and 5.1.2 may oscillate around a steady-state value. In order to account for this, an oscillation resistant-algorithm detailed in the supplementary material is employed.
The standard set of parameters employed is shown in Table 1, which will be used unless mentioned otherwise; these are nondimensional values and do not represent any specific material. The units they would have in dimensional form are listed together with them. As the concentration is measured in mole fraction, it is naturally dimensionless.
For equilibrium simulations, $D=100$ is used to speed up convergence. When performing grid convergence studies $W=3\Delta x^{0.4}$ is used as described in section 3.4 to ensure we converge to the sharp interface limit.
@@
parameter value units
$M_{\phi}$ 1 $\mathrm{m}^{4}\mathrm{/}\mathrm{(}\mathrm{J}\mathrm{s}\mathrm{)}$
$\gamma$ 1 $\mathrm{J}\mathrm{/}\mathrm{m}^{2}$
$W$ 2.5 $\mathrm{m}$
$\Delta x$ 1 $\mathrm{m}$
$D$ 1 $\mathrm{m}^{2}\mathrm{/}\mathrm{s}$
$k_{\alpha}=k_{\beta}$ 500 $\mathrm{J}\mathrm{/}\mathrm{m}^{3}$
$c_{0,\alpha}$ 0.02 -
$c_{0,\beta}$ 0.98 -
A short summary of the integrators used in this work together with how they are run is given here; a tabular overview is given in Table 2. The FEuler and SSP(5)2 schemes are run at fixed time steps, limited by their respectively stability limits. The SSP(10)4 scheme at tolerances of $1\text{\times}{10}^{-14}$ for all fields is used as a reference solution generator for testing convergence order for some problems. STS schemes of order 1 and 2 are run with both fixed time steps and adaptively. The time steps are based on being a multiple of the stable Euler time step up to a predetermined number. When integrating adaptively, we generally use a relative tolerance of $r=$1\text{\times}{10}^{-4}$\$. The absolute tolerance for the concentration is $a_{c}=$1\text{\times}{10}^{-4}$\$ unless mentioned otherwise; this represents about 1% of the smallest expected concentration value. As mentioned above, the phase-field tolerance will be varied between $1\text{\times}{10}^{-2}$ and $1\text{\times}{10}^{-4}$ and its influence mentioned where appropriate. When performing grid convergence studies, only the STS2 integrator with $a_{\phi}\in\{$1\text{\times}{10}^{-2}$,$1\text{\times}{10}^{-4}$\$\} will be used.
Table 2: Employed integrator configurations
integrator configurations
FEuler $\Delta t\in\{0.5,1\}\cdot\Delta t_{e}$
SSP(5)2 $\Delta t\in\{0.5,1,2,4\}\cdot\Delta t_{e}$
STS[12]
$\Delta t\in\{1,10,100,200\}\cdot\Delta t_{e}$
5 Results & discussion
In this section the geometry of the benchmarks together with their solution is presented. The first subsection will deal with equilibrium problems to address questions of energy stability, with the second one dealing with kinetic benchmarks to assess the computational efficiency of the schemes. The final subsection uses a geometry without analytical solution to show how the small-scale benchmark results translate into more practical applications. Performance is reported here in terms of RHS evaluations, as this measure is independent of the machine employed and hence eases future comparison with other time integrators. Qualitative comparison to implicit integrators can also be done with this as long as the resulting systems are solved with iterative methods as each iteration requires RHS evaluations; a more quantitative comparison needs to account for e.g. additional matrix-vector operations required by the solution algorithm as well as potential mass matrix inversions. The supplementary material includes data on the runtimes as well for the interested reader, though the conclusions do not change if performance is measured in runtime instead of RHS evaluations.
Finally, we generally output at defined target times exactly, with these being available for each configuration in the supplementary data. For fixed-step integrators this only manifests in an extra RHS evaluation for each output time not being an integer multiple of the time step width. For adaptive integrators, this generally reduces the step taken prior to reaching the output significantly and hence becomes less efficient than if a simple time crossing output strategy were adopted. Hence the performance results are slightly biased against adaptive integrators.
5.1 Equilibrium verification
An important part of the phase-field theory is that the derived PDEs represent gradient flows of an energy functional. This ensures that the evolution minimizes that functional. Schemes ensuring this are generally referred to as energy stableGomez2017 , with many classical schemes not being unconditionally so. In lieu of proving (conditional) energy stability for schemes employed herein, two cases with analytical equilibrium state are considered. Even if the schemes should temporarily have increasing energy (as shown), it is deemed acceptable if the analytical equilibrium state is the convergence point of a grid refinement study and that at a practical resolution, the scheme approximates the equilibrium state closely as well. Note that all schemes employed here contain at least one (scaled) forward Euler step and hence would at best be conditionally energy stable. We further note that even energy stable schemes might not lead to the correct kineticsXu2019 , so they should also be benchmarked before use.
5.1.1 Phase embedding
The first equilibrium case is that of one phase being embedded in another, with curvature providing a driving force for shrinkage which is counteracted by chemical driving forces. Here we solve both eq. 6 and eq. 17. Both phases are initialized with their flat equilibrium concentration values and thus there is a small driving force for diffusion in order to reach equilibrium in which there is a pressure difference, the Laplace pressure,
@@ 5.1.1 Phase embedding
[…]
@@
We first perform a grid convergence study with this geometry, resolving a domain of $128^{2}$ with an initial seed of size $r=32$. Equilibrium for this case is defined as a rate of change for the Laplace pressure $|\frac{(\Delta P(t)-\Delta P(t_{b}))}{t-t_{b}}|$ being smaller than $1\text{\times}{10}^{-14}$ for two times $t>t_{b}$ at least two diffusion times $128^{2}/D$ apart.
(a) errors in the Laplace pressure
(b) energy profile over time
5.1.2 Double triple junction
The second case is a triple junction geometry, now also probing variable parameters and multiple phases interacting. Again, both eq. 6 and eq. 17 are solved in this case. One phase $\alpha$ is placed on the grain boundary between two grains of the phase $\beta$, as sketched in fig. 2 together with the employed boundary conditions and geometric quantities, forming two triple junctions. The junctions will seek to equilibrate the surface tensions and hence evolve the structureBalluffi2005 . In contrast to Daubner2023 , we want to test the equilibrium together with a chemical driving force, and so do not require Dirichlet boundary conditions to prevent curvature minimization from eliminating all interfaces. We note that this choice generally leads to much longer simulation times, as the equilibrium is now linked to the long-range diffusion field. The evolution should be such that the phase $\alpha$ obtains an equilibrium shape defined by the dihedral angle
@@ 5.1.2 Double triple junction
with its momentary form being computable with the geometric parameters $L$, $S$ following Bradley1977 by assuming the shape is a vesica piscis:
@@
The phase $\alpha$’s distance field is given by eq. 44, with those of the two $\beta$ grains being $d_{\beta_{1}}=C_{x}-x$ and its complement respectively. In this case the sum of the phase-fields do not sum to 1 and hence the system eq. 41 is solved while fixing $\phi_{\alpha}$ and determining new values for both $\phi_{\beta}$.
Here we resolve a domain of size $192\times 96$ at the coarsest scale with $\Delta x=1$, with the long side being along the grain boundary to facilitate simulation of smaller dihedral angles at the same aspect ratio. The initial radius $r$ of the circular phase on the grain boundary is $32$. For simplicity the convergence study is limited to a single angle $\gamma_{\beta\beta}=1,\gamma_{\alpha\beta}=2\rightarrow\theta\approx$151\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, as results should be transferable to other angles except when approaching $\theta\to 0$ as complete dewetting is not captured by the model Daubner2023 . In addition, the coarsest resolution is also run with the FEuler integrator with a time step 0.9 times its stable time step. The remaining angles in the range $95\text{\,}\mathrm{\SIUnitSymbolDegree}180\text{\,}\mathrm{\SIUnitSymbolDegree}$ will use the practical resolution $\{\Delta x=1,W=2.5\}$ with $\gamma_{\alpha\beta}\in\{0.75,1,2,6\}$.
Equilibrium is defined as a simple approximation to the measurement described in section 5.1.2. The positions of the 0.5 lines are determined on the horizontal and vertical lines passing through the center of the domain, using only a linear interpolation in the respective coordinate directions, with their respective differences giving approximations to $L,S$. This allows an estimate of the current dihedral angle $\theta$, with equilibrium defined as a rate of change in angle by less than $1\text{\times}{10}^{-11}\text{\,}\mathrm{rad}$ over at least two diffusion times. The reported angle values are based on a more accurate marching squares reconstruction of the 0.5 contour lines of the phase $\alpha$, which is only done after the simulation is completed.
(a) errors in the dihedral angle
(b) energy profile over time
(c) time evolution of the angle
The results of this study are shown in fig. 3. Again, the grid convergence study in fig. 3(a) shows convergence to the sharp interface with little influence from $a_{\phi}$. At the coarsest resolution, a relative error of $1.4\%$ is obtained. In this case, the energy does not decrease monotonically beyond the first output frame as shown in the inset of fig. 3(b), though there is a long-term decrease of energy. As is visible from the graph, on the scale of the energy reduction it is hardly noticeable without zooming closely to the data. Note that the FEuler integrator at the coarsest resolution also exhibits a non-monotonic behaviour as shown in the top inset. We note though that energy instability was only apparent once the simulation was close to equilibrium; benchmarks not getting close to equilibrium are unlikely to probe energy stability properties of the integrator. If desired, one way of avoiding energy instability would be adding a second criterion for step rejection dependent on whether the energy decreased or not. If the energy increased, integration tolerances would be tightened and a new time step computed based on this until the energy no longer increases. This is based on the thought that error in energy is bounded by the change in the solution per step, with tightening tolerances reducing the change and hence reducing the energy error.
@@
5.2 Kinetic verification
Given that energy stability seems to play no practical role for equilibrium, we proceed with the kinetic benchmarks to determine the computational efficiency of the integrators.
5.2.1 Single grain growth
A single grain embedded within another grain will shrink due to excess energy contained in the interface, given that there are no other driving forces preventing this. Hence in this case we will use eq. 6 alone without concentration coupling, which means the sole driving force is the interfacial energy reduction. Given that curvature flow is the only active mechanism, one may write for the volume of the embedded grainBalluffi2005